OpenStudy (anonymous):

The midpoint of is (5, –11). The coordinates of one endpoint are U(3, 5). Find the coordinates of endpoint V.

OpenStudy (anonymous):

\[\frac{ 3+x }{ 2} = 5 \] 3+x=10 x=7 \[\frac{ 5+y }{ 2 }= - 11\] 5+y=-22 y=-27 Answer: (7,-27)

OpenStudy (anonymous):

That's what I got! Glad I am correct! Thanks!

OpenStudy (anonymous):

You are welcome! (: